OpenStudy (anonymous):

I'm solving exponential and logarithmic equations and I have no idea what I'm doing. The equation is 8^(2x)=32

OpenStudy (anonymous):

you write \[8=2^3\] and so \[\large 8^{2x}=(2^3)^{2x}=2^{6x}\]

OpenStudy (anonymous):

then since \(32=2^5\) you have \[\large 2^{6x}=2^5\] making \(6x=5\) and you can solve for \(x\)

OpenStudy (phi):

sat explained how to do it. The idea is IF you can write your equation to look like \[ 2^{6x} = 2^5\] then that means the exponents must be equal, i.e. 6x= 5 solving for x is "normal" algebra: divide both sides by 6 to find x= 5/6 So how do you change \[ 8^{2x} = 32 \] into that form ?

OpenStudy (phi):

In these simple problems, look for both numbers (the 8 and the 32) to be powers of the same base. 2 divides into 8 to get 2*4. and 4 is 2*2, so 8= 2*2*2 or 2^3 your hope is 32 is also a power of 2. and it is, 32= 2^5